calculate the boiling point of a solution containing 12g C6H12O6 dissolved in 200g of water H2O. (0.512oC - K6 of H2O)

1
Expert's answer
2021-05-31T07:44:56-0400

Molarity =0.243÷0.3599=0.676M


"\Delta H=(1)(0.676)(0.512)"


"\Delta H=0.35°C"